OpenStudy (anonymous):

Please help!!! Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (anonymous):

@hartnn @ganeshie8

hartnn (hartnn):

Square and add. basically we are using the identity \(\sin^2 a+\cos^2a=1\)

OpenStudy (anonymous):

x=4cos t cos t =x/4 similarly sin t = y/4 cos^2 t +sin^2 t =1 x2 /16 + y2/16 =1 x2 + y2 =16

hartnn (hartnn):

yes^

OpenStudy (anonymous):

wait so it should be y= 4 sin x/4 ?

OpenStudy (anonymous):

because it has to be the equation of y

hartnn (hartnn):

no no, first, square both the equations, what u get ?

OpenStudy (anonymous):

16?

hartnn (hartnn):

x = 4 cos t, -----> x^2 = 16 cos^2 t ----> (1) y = 4 sin t ------> y^2 = 16 sin^2 t ----->(2) got this ?

OpenStudy (anonymous):

okay

hartnn (hartnn):

now just add (1) and (2)

OpenStudy (anonymous):

okay I got it now lol :) thanks!

hartnn (hartnn):

welcome ^_^